# Modulo Sampling: Performance Guarantees in The Presence of Quantization

### v1

@@ II-B Signal Recovery for Modulo ADC With 1-Bit Folding Information
The signal recovery block utilizes the information about $\mathcal{S}$ to recover the modulo residue $z[n]$. The underlying reconstruction algorithm is based on [16, 32]. Since $f(t)\in L^{1}(\mathbb{R})$, $\lim_{|t|\rightarrow\infty}f(t)=0$ by the Reimann-Lebesgue lemma [33, Chapter 12]. Consequently, for any $\lambda^{\prime}>0$, there exist integers $n_{0}$ and $n_{1}$ ($n_{0}<n_{1}$) such that $|f(nT_{\mathrm{s}})|<\lambda$ for all $n_{0}>n$ and $n_{1}<n$. the residual samples $z[n]$ can be treated as a finite-duration discrete-time signal of length $N=n_{1}-n_{0}+1$. We define a $K\times 1$ vector $\mathbf{y}=[y_{0},\cdots,y_{K-1}]^{T}$, where
$y_{k}=$ $\frac{1}{\sqrt{N}}\sum_{n=n_{0}}^{n_{1}}\left[f_{\lambda^{\prime},\mathrm{q}}[n]-f_{\lambda^{\prime},\mathrm{q}}[n-1]\right]e^{-j\frac{2\pi nk}{N}}$
$=$ $\frac{1}{\sqrt{N}}\sum_{n=n_{0}}^{n_{1}}\left[\Delta z[n]+\Delta\epsilon[n]\right]e^{-j\frac{2\pi nk}{N}}$
$=$ $\frac{1}{\sqrt{N}}\Big{(}\sum_{n\in\mathcal{S}}\left[\Delta z[n]+\Delta\epsilon[n]\right]e^{-j\frac{2\pi nk}{N}}$
$\quad+\sum_{n=n_{0},n\notin\mathcal{S}}^{n_{1}}\Delta\epsilon[n]e^{-j\frac{2\pi nk}{N}}\Big{)}.$
[…]
Let $\mathbf{V}\in\mathbb{R}^{K\times N}$ with elements $\mathbf{V}_{k,n}=e^{-j\frac{2\pi kn}{N}}$, where $\frac{2\pi k}{N}$ is the $k$-th discrete frequency in $(\rho\pi,2\pi-\rho\pi)$. The matrix $\mathbf{V}_{\mathcal{S}}\in\mathbb{R}^{K\times|\mathcal{S}|}$ contains the $|\mathcal{S}|$ columns of $\mathbf{V}$ corresponding to the indices in $\mathcal{S}$. An estimate of the non-zero elements of the first-order difference vector, denoted ${\Delta}\mathbf{\hat{z}}_{\mathcal{S}}$, can be formed by
@@
The estimate of the vector ${\Delta}\mathbf{z}$, denoted ${\Delta}\mathbf{\hat{z}}$, is obtained by combining ${\Delta}\mathbf{\hat{z}}_{\mathcal{S}}$ and ${\Delta}\mathbf{\hat{z}}_{\mathcal{S}^{c}}$. The elements of ${\Delta}\mathbf{\hat{z}}$ are rounded to the nearest $2\lambda^{\prime}\mathbb{Z}$. Then a cumulative sum is applied to the resulting vector to obtain the estimate of the modulo residue, denoted $\mathbf{\hat{z}}=({\Delta}\mathbf{\hat{z}})_{\Sigma}$. Let $\hat{z}[n]$ be the discrete-time representation of the vector $\mathbf{\hat{z}}$. We subtract $\hat{z}[n]$ from $f_{\lambda^{\prime},\mathrm{q}}[n]$ to remove the modulo residue from the quantized modulo observations. Finally, we apply an ideal digital lowpass filter, $\mathrm{LPF}\{\cdot\}$, with passband region $\left(-\frac{\omega_{m}}{2},\;+\frac{\omega_{m}}{2}\right)$. The recovered signal can be written as
@@
$=$ $f[n]+\mathrm{LPF}\left\{z[n]-\hat{z}[n]\right\}+\mathrm{LPF}\{\epsilon[n]\}.$ (5)
Hence, the recovered signal is composed of the desired samples, filtered modulo residue estimation error, and filtered quantization noise.
@@ II-C Modulo ADC Without 1-Bit Folding Information
In a typical modulo ADC system, the 1-bit folding information signal may not be available for the reconstruction method. Thus, it is also of interest to develop a recovery algorithm for the modulo ADC shown in Fig. 2b. Since the 1-bit folding information tells the exact locations of nonzero elements of $\Delta z$, one approach to facilitate the signal reconstruction is to perform support recovery of $\Delta z[n]$ and then run the signal reconstruction algorithm described in Section II-B, with $c[n]$ being replaced by $\hat{c}[n]$, the recovered support of $\Delta z[n]$. Note that $\Delta z[n]$ is a sparse signal [15]. Thus, we can utilize existing support recovery algorithms for sparse signals and their corresponding theoretical guarantees.
For the support recovery of $\Delta z[n]$, we consider the greedy orthogonal matching pursuit (OMP) algorithm [34, 35], which uses a stopping rule based on the $\ell_{\infty}$-norm of the bounded noise [36]. More precisely, $\hat{c}[n]$ is obtained as follows:
@@
    Otherwise, set iteration counter to $t=t+1$ and return to Step 2.
@@
Our aim is to derive MSE performance guarantees for the two modulo ADC systems presented when the proposed recovery algorithms are used. Additionally, we also want to identify suitable settings for which the modulo ADC has superior performance compared to a conventional ADC without modulo. Furthermore, we extend this analysis to a critical scenario involving the simultaneous acquisition of weak and strong signals occupying different frequency bands in a BL signal. The following sections present these performance guarantees in detail.
III Recovery Guarantees
We quantify the recovery performance of the modulo ADC using the MSE criterion. In general, the MSE can be expressed as a sum of two errors: (1) a truncation error due to neglecting the signal values outside $n<n_{0}$ and $n>n_{1}$, and (2) a reconstruction error due to the modulo residue estimation error and finite quantization. In this paper, we focus on the latter since we can set $n_{0}$ and $n_{1}$ so that $f[n]\approx 0$ outside the discrete-time interval $[n_{0},n_{1}]$. Consequently, the truncation error can be set to a negligible value. Henceforth, we refer to the reconstruction error whenever MSE is mentioned.
@@ III-A MSE guarantee for Modulo ADC With 1-Bit Folding Information
Suppose $\mathrm{OF}>3$. Then $\mathbf{V}_{\mathcal{S}}$ has full column rank if the modulo threshold $\lambda^{\prime}\geq\frac{\|f(t)\|_{\infty}}{\mathrm{OF}-2}$.
@@
For a given $f(t)$, it actually suffices that $\mathrm{OF}\geq\frac{N}{N-|\mathcal{S}|}$ to make $\mathbf{V}_{\mathcal{S}}$ full column rank. However, such condition depends on the specific $f(t)$. By using the upper bound on $|\mathcal{S}|$ established in [15], we eliminate the dependence on the specific $f(t)$ and find an oversampling factor requirement that holds for any $f(t)$. However, the bound on $|\mathcal{S}|$ is not tight in general as demonstrated in [15, Table 1].
@@
Suppose $b>3$ and $\mathrm{OF}>3$. Then the MSE of the reconstruction procedure for the modulo ADC with extra 1-bit folding information is
$\mathbb{E}[(f[n]-\hat{f}[n])^{2}]=\frac{1}{\mathrm{OF}(2^{b}-2)^{2}}\cdot\left(\frac{\|f(t)\|_{\infty}}{\mathrm{OF}-2}\right)^{2}$ (6)
if we set $\lambda^{\prime}=\frac{\|f(t)\|_{\infty}}{\mathrm{OF}-2}$.
@@
The intuition of the proof is to set a sufficiently large number of quantization bits such that the induced quantization noise is not strong enough to put the recovered first-order difference of the residue outside $(2\lambda^{\prime}p-\lambda^{\prime},2\lambda^{\prime}p+\lambda^{\prime})$ if $z[n]=2\lambda^{\prime}p$ for some $p\in\mathbb{Z}$. In this case, the rounding operation on the elements of ${\Delta}\hat{z}$ is perfect, i.e. $\hat{z}[n]=z[n]$, and the only impairment in $\hat{f}[n]$ is the (lowpass-filtered) quantization noise.
We now compare our derived MSE for the modulo ADC in Fig. 2a with that of the conventional ADC without modulo sampling. To get the behavioral model of a conventional ADC, we simply replace $\mathcal{M}_{\lambda^{\prime}}(\cdot)$ in Fig. 2a with a regular sampler with sampling period $T_{\mathrm{s}}$ and then remove $c[n]$. To obtain analytical results for conventional ADCs, we also use the non-subtractive dithered quantization framework. With triangle dither $d[n]\in\left(-\frac{2\lambda}{2^{b}},\;+\frac{2\lambda}{2^{b}}\right]$, we must set $\lambda$ of the conventional ADC as $\lambda=\frac{2^{b}}{2^{b}-2}\|f(t)\|_{\infty}$ to prevent overloading. Consequently, the quantization noise is white and has a mean squared value of $\mathbb{E}\left[\epsilon^{2}[n]\right]=\frac{\lambda^{2}}{2^{2b}}$. With an oversampling factor of $\mathrm{OF}$, the desired signal occupies only $\frac{1}{\mathrm{OF}}$ of $(-\pi,+\pi)$ band. A lowpass filter with passband region $(-\rho\pi,+\rho\pi)$ can be used to filter out the quantization noise in the out-of-band region. Thus, the MSE is solely attributed to the filtered quantization noise and can be expressed as
@@
We also compare our results with other theoretical guarantees derived for modulo sampling. The impact of bounded noise (e.g., quantization) has been analyzed in [9, Theorem 3] when the recovery algorithm is based on higher-order differences (HOD). Their analysis guarantees the recovery of the samples up to an unknown additive constant, i.e., $\hat{f}^{\mathrm{(HOD)}}[n]=f[n]+\epsilon[n]+2\lambda^{\prime}p$ for some unknown $p\in\mathbb{Z}$. By ignoring the unknown additive constant and applying an appropriate noise filtering, the MSE of the HoD should be identical to that of our proposed recovery algorithm. However, under this bounded noise setting, the sufficient condition to achieve this performance guarantee is that the sampling rate should be at least $2^{\alpha}\pi e$ times the Nyquist rate, where $\alpha\in\mathbb{N}$ is a noise-dependent parameter. In contrast, our analytical result does not have an unknown additive constant and the exact MSE can be derived at any $\mathrm{OF}>3$.

### v2

@@ II-B Signal Recovery for Modulo ADC With 1-Bit Folding Information
The signal recovery block utilizes the information about $\mathcal{S}$ to recover the modulo residue $z[n]$ in the frequency domain. Let $\Delta f_{\lambda^{\prime},\mathrm{q}}[n]=\Delta f[n]+\Delta z[n]+\Delta\epsilon[n]$ be the first-order difference of $f_{\lambda^{\prime},\mathrm{q}}[n]$. The discrete-time Fourier Transform (DTFT) of $\Delta f_{\lambda^{\prime},\mathrm{q}}[n]$ for $\Omega\in(\rho\pi,2\pi-\rho\pi)$ can be expressed as
$F_{\Delta,\mathrm{OOB}}(e^{j\Omega})$ $=\sum_{n=-\infty}^{\infty}\Delta f_{\lambda^{\prime},\mathrm{q}}[n]e^{-j\Omega n}$
$=\sum_{n=-\infty}^{\infty}\left[\Delta z[n]+\Delta\epsilon[n]\right]e^{-j\Omega n}.$ (4)
The subscript $\mathrm{OOB}$ is used to denote that we are only interested in out-of-band (OOB) frequencies $\Omega\in(\rho\pi,2\pi-\rho\pi)$. The second line follows since the signal $f[n]$ is within $(-\rho\pi,\;+\rho\pi)$. As such, $\Delta f[n]$ has no contribution to $F_{\Delta,\mathrm{OOB}}(e^{j\Omega})$. Because $f(t)\in L^{1}(\mathbb{R})$, we have $\lim_{|t|\rightarrow\infty}f(t)=0$ by the Reimann-Lebesgue lemma [42, Chapter 12]. Consequently, for any $\lambda^{\prime}>0$, there exist integers $n_{0}$ and $n_{1}$ ($n_{0}<n_{1}$) such that $|f(nT_{\mathrm{s}})|<\lambda$ for all $n_{0}>n$ and $n_{1}\leq n$. The first-order difference of the residual samples, $\Delta z[n]$, can be treated as a finite-duration discrete-time signal of length $N=n_{1}-n_{0}$. Equation (II-B) can be written as
$F_{\Delta,\mathrm{OOB}}(e^{j\Omega})=$ $\sum_{n=n_{0}}^{n_{1}-1}\left[\Delta z[n]+\Delta\epsilon[n]\right]e^{-j\Omega n}$
$+\sum_{n\notin\{n_{0},\cdots,n_{1}-1\}}\Delta\epsilon[n]e^{-j\Omega n}.$ (5)
We approximate $F_{\Delta,\mathrm{OOB}}(e^{j\Omega})$ by setting $f[n]=0$ for $n<n_{0}$ and $n_{1}$ and defining $y_{k}$ to be
$y_{k}=$ $\frac{1}{\sqrt{N}}\sum_{n=n_{0}}^{n_{1}-1}\left[\Delta z[n]+\Delta\epsilon[n]\right]e^{-j\frac{2\pi k(n-n_{0})}{N}}$
$=$ $\frac{1}{\sqrt{N}}\Big{(}\sum_{n=n_{0},n\in\mathcal{S}}^{n_{1}-1}\left[\Delta z[n]+\Delta\epsilon[n]\right]e^{-j\frac{2\pi k(n-n_{0})}{N}}$
$\quad+\sum_{n=n_{0},n\notin\mathcal{S}}^{n_{1}-1}\Delta\epsilon[n]e^{-j\frac{2\pi k(n-n_{0})}{N}}\Big{)}.$ (6)
Here, $\frac{2\pi k}{N}$ is one of the $K<N$ discrete frequencies inside $(\rho\pi,\;2\pi-\rho\pi)$. Effectively, $y_{k}$ is the normalized DFT of the truncated $\Delta f_{\lambda^{\prime},\mathrm{q}}[n]$ in the OOB region. The second line follows from putting all terms with $n\in\mathcal{S}$ in the first summation and putting the remainder in the second summation. Notice that the residual samples only appear in the first summation. We vectorize $\Delta z[n]$ by forming the vector ${\Delta}\mathbf{z}=[\Delta z[n_{0}],\cdots,\;\Delta z[n_{1}-1]]^{T}$. We also define the $|\mathcal{S}|\times 1$ vector ${\Delta}\mathbf{z}_{\mathcal{S}}$ which contains the nonzero elements of ${\Delta}\mathbf{z}$ at indices found in the set $\mathcal{S}$ while the $(N-|\mathcal{S}|)\times 1$ vector ${\Delta}\mathbf{z}_{\mathcal{S}^{c}}$ is an all-zero vector.
Let $\mathbf{V}\in\mathbb{R}^{K\times N}$ with elements $\mathbf{V}_{k,n}=e^{-j\frac{2\pi k(n-n_{0})}{N}}$, where $\frac{2\pi k}{N}$ is the $k$-th discrete frequency in $(\rho\pi,2\pi-\rho\pi)$. The matrix $\mathbf{V}_{\mathcal{S}}\in\mathbb{R}^{K\times|\mathcal{S}|}$ contains the $|\mathcal{S}|$ columns of $\mathbf{V}$ corresponding to the indices in $\mathcal{S}$. An estimate of the non-zero elements of the first-order difference vector, denoted ${\Delta}\mathbf{\hat{z}}_{\mathcal{S}}$, can be formed by
@@
The estimate of the vector ${\Delta}\mathbf{z}$, denoted $^{\Delta}\mathbf{\hat{z}}$, is obtained by combining ${\Delta}\mathbf{\hat{z}}_{\mathcal{S}}$ and ${\Delta}\mathbf{\hat{z}}_{\mathcal{S}^{c}}$. The elements of $^{\Delta}\mathbf{\hat{z}}$ are rounded to the nearest $2\lambda^{\prime}\mathbb{Z}$. Then, a cumulative sum is applied to the resulting vector to obtain the estimate of the modulo residue, denoted $\mathbf{\hat{z}}=({\Delta}\mathbf{\hat{z}})_{\Sigma}$. Let $\hat{z}[n]$ be the discrete-time representation of the vector $\mathbf{\hat{z}}$. We subtract $\hat{z}[n]$ from $f_{\lambda^{\prime},\mathrm{q}}[n]$ to remove the modulo residue from the quantized modulo observations. Finally, we apply an ideal digital lowpass filter, $\mathrm{LPF}\{\cdot\}$, with passband region $\left(-\frac{\omega_{m}}{2},\;+\frac{\omega_{m}}{2}\right)$. The recovered signal can be written as
@@
$=$ $f[n]+\mathrm{LPF}\left\{z[n]-\hat{z}[n]\right\}+\mathrm{LPF}\{\epsilon[n]\},$ (8)
for $n\in\{n_{0},\cdots,n_{1}-1\}$. Hence, the recovered signal is composed of the desired samples, filtered modulo residue estimation error, and filtered quantization noise.
@@ II-C Modulo ADC Without 1-Bit Folding Information
In a typical modulo ADC system, the 1-bit folding information signal may not be available for the reconstruction method. Thus, it is also of interest to develop a recovery algorithm for the modulo ADC shown in Fig. 2b. Since the 1-bit folding information tells the exact locations of nonzero elements of $\Delta z$, one approach to facilitate the signal reconstruction is to perform support recovery of $\Delta z[n]$ and then run the signal reconstruction algorithm described in Section II-B, with $c[n]$ being replaced by $\hat{c}[n]$, the recovered support of $\Delta z[n]$. Note that $\Delta z[n]$ is a sparse signal [20, Section III]. Thus, we can utilize existing support recovery algorithms for sparse signals and their corresponding theoretical guarantees.
For the support recovery of $\Delta z[n]$, we consider the greedy orthogonal matching pursuit (OMP) algorithm [43, 44, 45], which uses a stopping rule based on the $\ell_{\infty}$-norm of the bounded noise [46]. More precisely, $\hat{c}[n]$ is obtained as follows:
@@
    Otherwise, set the iteration counter to $t=t+1$ and return to Step 2.
@@
Our aim is to derive MSE performance guarantees for the two modulo ADC systems presented when the proposed recovery algorithms are used. The following section establishes these performance guarantees.
III Recovery Guarantees and Computational Complexity
We quantify the recovery performance of the modulo ADC using the MSE criterion, which can be mathematically expressed as
$\mathrm{MSE}=$ $\frac{1}{N}\sum_{n=n_{0}}^{n_{1}-1}\mathbb{E}\left\{\big{|}\hat{f}[n]-f[n]\big{|}^{2}\right\}$
$+\sum_{n\notin\{n_{0},\cdots,n_{1}-1\}}\mathbb{E}\{|f[n]|^{2}\},$ (9)
where $N=n_{1}-n_{0}$, as defined earlier. The second summation accounts for the truncation error, denoted $\mathrm{MSE}_{t}$, due to neglecting the signal values outside $\{n_{0},\cdots,n_{1}-1\}$ while the first summation accounts for the reconstruction error, denoted $\mathrm{MSE}_{r}$, due to the modulo residue estimation error and finite quantization. In this paper, we focus on the latter since we can set $n_{0}$ and $n_{1}$ so that $f[n]\approx 0$ for $n\notin\{n_{0},\cdots,n_{1}-1\}$. Consequently, the truncation error can be set to a negligible value. Henceforth, we refer to the reconstruction error whenever MSE is mentioned. The reconstruction error can be expressed as
$\mathrm{MSE}_{r}=$ $\frac{1}{N}\sum_{n=n_{0}}^{n_{1}-1}\mathbb{E}\left\{\big{|}\hat{f}[n]-f[n]\big{|}^{2}\right\}$
$=$ $\frac{1}{N}\sum_{n=n_{0}}^{n_{1}-1}\bigg{|}\mathrm{LPF}\left\{z[n]-\hat{z}[n]\right\}\bigg{|}^{2}$
$+2\mathbb{E}\left\{\mathrm{LPF}\left\{\epsilon[n]\right\}\right\}\mathrm{LPF}\left\{z[n]-\hat{z}[n]\right\}$
$+\mathbb{E}\left\{\big{|}\mathrm{LPF}\left\{\epsilon[n]\right\}\big{|}^{2}\right\}$
$=$ $\frac{1}{N}\sum_{n=n_{0}}^{n_{1}-1}\bigg{|}\mathrm{LPF}\left\{z[n]-\hat{z}[n]\right\}\bigg{|}^{2}+\mathbb{E}\left\{\big{|}\mathrm{LPF}\left\{\epsilon[n]\right\}\big{|}^{2}\right\}.$ (10)
The second line follows from expanding the expression in the first line. Since $\epsilon[n]$ has zero mean, the second term of the second line can be dropped. The first term of the third line comes from the lowpass-filtered modulo residue estimation error while the second term of the third line comes from the in-band quantization noise. Our objective in this section is to derive sufficient conditions for $\mathrm{OF}$ and $b$ so that in-band quantization noise is the only source of reconstruction error for the modulo ADC with 1-bit folding information and modulo ADC without 1-bit folding information.
@@ III-A MSE guarantee for Modulo ADC With 1-Bit Folding Information
Suppose $\mathrm{OF}>3$ and $n_{0}$$n_{1}$ are chosen so that truncation error is negligible. Then $\mathbf{V}_{\mathcal{S}}$ has full column rank if the modulo threshold $\lambda^{\prime}\geq\frac{\|f(t)\|_{\infty}}{\mathrm{OF}-2}$.
@@
The intuition of the proof is to relate the oversampling factor $\mathrm{OF}$ to the dimensions of the sensing matrix $\mathbf{V}_{\mathcal{S}}$. Specifically, increasing $\mathrm{OF}$ has two key effects: (1) it increases the number of discrete frequencies in the out-of-band (OOB) region, thereby increasing the number of rows in $\mathbf{V}_{\mathcal{S}}$; and (2) it reduces the number of folding events in a length-$N$ signal, which corresponds to a decrease in the number of columns of $\mathbf{V}_{\mathcal{S}}$. The proof identifies a regime where the number of rows $K$ exceeds or matches the number of columns $|\mathcal{S}|$, ensuring that the sensing matrix is sufficiently tall to enable stable recovery under appropriate conditions.
For a given $f(t)$, it actually suffices that $\mathrm{OF}\geq\frac{N}{N-|\mathcal{S}|}$ to make $\mathbf{V}_{\mathcal{S}}$ full column rank. However, such a condition depends on the specific $f(t)$. By using the upper bound on $|\mathcal{S}|$ established in [20], we eliminate the dependence on the specific $f(t)$ and find an oversampling factor requirement that holds for any $f(t)$. However, the bound on $|\mathcal{S}|$ is not tight in general as demonstrated in [20, Table 1].
@@
Suppose $b>3$, $\mathrm{OF}>3$, and $n_{0}$$n_{1}$ are chosen so that truncation error is negligible. Then the $\mathrm{MSE}_{r}$ of the reconstruction procedure for the modulo ADC with extra 1-bit folding information is
$\mathrm{MSE}_{r}=\frac{c^{2}}{\mathrm{OF}(2^{b}-2)^{2}}\cdot\left(\frac{\|f(t)\|_{\infty}}{\mathrm{OF}-2}\right)^{2}$ (11)
if we set $\lambda^{\prime}=\frac{c\cdot\|f(t)\|_{\infty}}{\mathrm{OF}-2}$ for some positive constant $c\in[1,\mathrm{OF}-2]$.
@@
The intuition of the proof is to set a sufficiently large number of quantization bits such that the induced quantization noise is not strong enough to put the recovered first-order difference of the residue outside $(2\lambda^{\prime}p-\lambda^{\prime},2\lambda^{\prime}p+\lambda^{\prime})$ if $z[n]=2\lambda^{\prime}p$ for some $p\in\mathbb{Z}$. In this case, the rounding operation on the elements of $^{\Delta}\hat{z}$ is perfect, i.e., $\hat{z}[n]=z[n]$, and the only impairment in $\hat{f}[n]$ is the (lowpass-filtered) quantization noise.
We now compare our derived MSE for the modulo ADC in Fig. 2a with that of the conventional ADC without modulo sampling. To get the behavioral model of a conventional ADC, we simply replace $\mathcal{M}_{\lambda^{\prime}}(\cdot)$ in Fig. 2a with a regular sampler with sampling period $T_{\mathrm{s}}$ and then remove $c[n]$. To obtain analytical results for conventional ADCs, we also use the non-subtractive dithered quantization framework. With triangle dither $d[n]\in\left(-\frac{2\lambda}{2^{b}},\;+\frac{2\lambda}{2^{b}}\right]$, we must set $\lambda$ of the conventional ADC as $\lambda=\frac{2^{b}}{2^{b}-2}\|f(t)\|_{\infty}$ to prevent overloading. Consequently, the quantization noise is white and has a mean squared value of $\mathbb{E}\left[\epsilon^{2}[n]\right]=\frac{\lambda^{2}}{2^{2b}}$. With an oversampling factor of $\mathrm{OF}$, the desired signal occupies only $\frac{1}{\mathrm{OF}}$ of the whole DTFT spectrum. A lowpass filter with passband region $(-\rho\pi,+\rho\pi)$ can be used to filter out the quantization noise in the out-of-band region. Thus, the MSE is solely attributed to the filtered quantization noise and can be expressed as
@@
We also compare our results with other theoretical guarantees derived for modulo sampling. The impact of bounded noise (e.g., quantization) has been analyzed in [11, Theorem 3] when the recovery algorithm is based on HoD. Their analysis guarantees the recovery of the samples up to an unknown additive constant, i.e., $\hat{f}^{\mathrm{(HOD)}}[n]=f[n]+\epsilon[n]+2\lambda^{\prime}p$ for some unknown $p\in\mathbb{Z}$. By ignoring the unknown additive constant and applying an appropriate noise filtering, the MSE of the HoD should be identical to that of our proposed recovery algorithm. However, under this bounded noise setting, the sufficient condition to achieve this performance guarantee is that the sampling rate should be at least $2^{\alpha}\pi e$ times the Nyquist rate, where $\alpha\in\mathbb{N}$ is a noise-dependent parameter. In contrast, our analytical result does not have an unknown additive constant and the exact MSE can be derived at any $\mathrm{OF}>3$. A prediction-based method introduced in [18] demonstrated that the oversampling factor ($\mathrm{OF}$) can be reduced arbitrarily close to 1 by significantly increasing the length of the prediction filter. However, as noted in [18, Section III], the method becomes unreliable when quantization noise is present. In contrast, our approach remains effective even when using a finite number of quantization bits.
We would also like to mention key differences between our proposed approach and the UNO framework [47]. Unlike UNO, which relies on a large number of 1-bit measurements per sample using varying thresholds, our method processes a single quantized measurement per folded sample. Furthermore, our reconstruction guarantees are deterministic and hold for all finite-energy bandlimited signals, while UNO provides high-probability bounds. Lastly, our oversampling requirement is significantly lower, with a sufficient condition of $\mathrm{OF}\geq 3$ compared to $\mathrm{OF}\geq 2^{h}\pi e$, where $h\in\mathbb{N}$, required in [47]. We also observe that the condition $\mathrm{OF}>3$ is similar to the oversampling requirement established in [48] for the unique recovery of periodic bandlimited signals under a modulo-DFT sensing framework. Unlike our model, their approach applies the DFT to the sampled signal before the modulo operation—a structure that may be impractical to implement in real-world systems. Finally, we also note that sliding DFT approach of our proposed algorithm has been introduced in [19]. The recovery guarantees established in [19] adopted the recovery guarantees in the conference version of our work [41] but incorporated the impact of spectral leakage in the algorithm performance. In contrast, we assumed that spectral leakage can be neglected in our work if $n_{0}$ and $n_{1}$ are chosen appropriately.